# Identification of three new ULXs in NGC 4631 and NGC 1097: evidence for stellar-mass black holes

### v1

@@ Abstract
Context. Recent observations of galaxies continue to reveal new ultraluminous X-ray sources (ULXs), increasing their known population. This growing sample provides improved statistics essential for advancing our understanding of ULX properties and their nature.
Aims. Our aim is to study the source populations using published Chandra, XMM-Newton, and Swift/XRT observations to identify new ULXs. In particular, we focus on investigating their X-ray and optical properties to constrain the nature of the compact objects and donor stars.
Methods. We analyzed archival X-ray observations spanning 2000–2025 for NGC 4631 and for NGC 1097 to study their ULX populations. Spectral fitting was performed for sources with sufficient counts using absorbed power-law and diskbb models to determine their physical properties. Variability analyses, including hardness-intensity diagrams and light curves, were conducted to assess both short- and long-term variability. Additionally, optical color-magnitude diagrams (CMDs) and near-infrared (NIR) spectral energy distributions (SEDs) were used to identify and characterize possible donor stars.
Results. In NGC 4631, we identified two new transient ULXs, X-6 and X-7, whose X-ray count rates vary by more than an order of magnitude. The relation obtained from the diskbb fits provides strong evidence that NGC 4631 X-6 is powered by a stellar-mass black hole accreting via a standard disk. The optical sources within the X-6 and X-7 X-ray error circles are considered candidate optical counterparts, suggesting that these systems are candidate HMXBs. In NGC 1097, we report the discovery of a new transient ULX, designated ULX-3, exhibiting X-ray luminosity (LX) variations by a factor of about 30. For ULX-3, we detected spectral state transitions, which may indicate a compact object consistent with either a stellar-mass black hole (BH) or a neutron star (NS). Moreover, we identified a unique optical and NIR counterpart. While the optical emission is variable, the NIR emission remains stable. The NIR counterpart of ULX-3 shows an SED consistent with a blackbody temperature of K, which lies within the expected range of red supergiants (RSGs) with a stellar radius of .
Conclusions.
@@
Ultraluminous X-ray sources (ULXs) are the most extreme subclass of X-ray binaries (XRBs), off-nuclear point sources exhibiting ¿ 2 erg s-1 (Kaaret et al., 2017; Fabrika et al., 2021; King et al., 2023; Pinto and Walton, 2023). To account for their ultra luminosities, recent studies increasingly support the view that the majority of ULXs are powered by stellar-mass compact objects undergoing super-Eddington accretion (King et al. (2023) and references therein), rather than by sub-Eddington accretion onto intermediate-mass black holes (Heinzeller et al., 2006; Brightman et al., 2019; Greene, 2022). Within this framework, the accreting compact object is generally inferred to be either a BH or an NS.
@@
So far, coherent pulsations have been detected in several ULXs, confirming that they are pulsating NS ULXs (PULXs), with the pulsations identified primarily through observations by XMM-Newton and NuSTAR (Bachetti et al., 2014; Motch et al., 2014; Israel et al., 2017; Carpano et al., 2018). Moreover, the observation of cyclotron resonance scattering features in the X-ray spectrum of an ULX implies the presence of an NS candidate (Mushtukov et al., 2017; Pinto et al., 2020). In addition, several ULXs, including some identified as PULXs, show transient behavior, whereas most remain persistently bright in the X-ray band over long timescales. Repeated observations of nearby galaxies with Chandra and XMM-Newton have contributed to a growing number of ULXs, particularly those exhibiting transient behavior (van Haaften et al., 2019; Earnshaw et al., 2020; Walton et al., 2021; Robba et al., 2022; Roberts et al., 2023; Brightman et al., 2023; Allak, 2023).
@@
Several ULXs have also shown clear evidence of recurrent spectral state transitions, evolving from hard to soft regimes over time, as revealed by high-cadence and long-term studies. In particular, Holmberg II X1 and NGC 5204 X1 show a cyclical evolution through the states of hard UL (HUL), soft UL (SUL), and supersoft UL (SSUL), probably driven by variations in accretion rate and changes in funnel geometry caused by radiation‑driven winds (Gúrpide et al., 2021b; Mondal et al., 2021). The recent study of Majumder et al. (2025) analyzed the long-term evolution of several ULXs with candidate BH accretors using XMM–Newton, revealing diverse spectrotemporal behaviors and correlations between disc luminosity (Ldisc) and inner disc temperature (Tcol). They argued that these correlations are indicative of possible spectral states or transitions between standard thin disks and slim disk configurations at different luminosities, supporting a two-component supercritical accretion framework regulated by the disc, corona, and radiation-driven winds. Numerous studies have aimed to identify the optical counterparts of ULXs and clarify the origin of their emission (Tao et al., 2011; Soria et al., 2012; Gladstone et al., 2013; Vinokurov et al., 2018; Allak et al., 2022; Guo et al., 2023; Allak and Akyuz, 2025). However, determining the precise contributions remains challenging, as the emission may originate from the donor star’s photosphere, the accretion disc, or a combination of both (Tao et al., 2011; Fabrika et al., 2015; Allak et al., 2022). In rare cases, spectral signatures such as stellar absorption lines allow clear donor identification, as seen in NGC 7793 P13, where a B9Ia supergiant was confirmed (Motch et al., 2014). Optical variability has also been observed in several ULXs, such as ULX-4 and ULX-8 in M51 and NGC 1313 X-2, suggesting accretion disc-dominated emission in these systems (Liu et al., 2009; Allak, 2022). Optical counterparts of ULXs are generally faint (mV ¿ 20 mag) and frequently obscured by surrounding stellar populations or interstellar material, which complicates their detection and interpretation. In contrast, infrared (IR) observations, particularly in the NIR range, are less affected by dust extinction and offer valuable insights into the nature of donor stars and the circumstellar environments in ULXs. Several ULXs have been reported to have NIR magnitudes consistent with red supergiants (RSGs) Heida et al., 2014; Lau et al., 2019; Dudik et al., 2016, and spectroscopic analyses have suggested RSG-like donors in some cases (Heida et al., 2016). Mid-IR emission has also been attributed to either circumstellar dust or jets (Dudik et al., 2016; Lau et al., 2019; Allak, 2023). NGC 4631, also known as the Whale Galaxy, is a barred spiral galaxy exhibiting starburst activity. It is located in the constellation Canes Venatici, approximately 7.6 Mpc away (1′′ = 34 pc) (Soria and Ghosh, 2009). Several studies have explored the ULX population in NGC 4631, revealing a range of accretion processes. Carpano et al. (2007) reported the discovery of a supersoft ULX that exhibits a 4-hour periodic modulation based on XMM-Newton data. This supersoft ULX is unlikely to be explained by the standard white dwarf scenario without extreme beaming. It has been suggested that it may instead be a stellar-mass BH accreting at super-Eddington rates, with determined modulation resulting from either a companion eclipse or a warped accretion disc.
@@
Furthermore, using archival data from XMM-Newton, Chandra, and ROSAT, Soria and Ghosh (2009) reanalyzed the five brightest X-ray sources in NGC 4631, identifying four (X1, X2, X4, and X5) as ULXs. Among these, X1—previously described as a supersoft source by Carpano et al. (2007) has been proposed to represent an extreme subclass of supersoft sources, powered by transient super-Eddington outbursts driven by non-steady nuclear burning on the surface of a massive white dwarf. The other sources displayed spectral properties consistent with stellar-mass BHs accreting at super-Eddington rates. Moreover, recently, Ducci et al. (2025) reported the discovery of a new transient pulsar ULX in NGC 4631, designated X–8, with a spin period of and a very high spin-up rate, . The optical counterparts of ULXs in NGC 4631 were studied by Guo et al. (2023) through a broadband and narrowband imaging campaign with CFHT/MegaCam. They sought to detect bubble structures around the X-ray sources and to examine their accretion properties. The stellar surroundings of the sources were analyzed via extinction-corrected color–magnitude diagrams and isochrone fitting. Their study revealed a highly asymmetric bubble nebula around X4, displaying distinct morphologies in H and [O III] emission.
@@
NGC 1097 (D = 16.8 Mpc, 1′′ = 82 pc; Liu 2011) is typically classified as a barred spiral (SBb) galaxy. It is also classified as a Seyfert 2 galaxy, which hosts an active galactic nucleus (AGN) powered by a central supermassive BH. Consequently, most studies of this galaxy have focused on constructing the SED of its nucleus using X-ray and multiwavelength observations (Nemmen et al., 2006). Chandra data further revealed two ULX sources, X2 and X3, in the galaxy, with model-estimated luminosities of erg s-1 and erg s-1, respectively (Liu, 2011).
@@
In this study, we investigate the nature of two newly identified ULX sources in NGC 4631, designated X-6 and X-7, and one ULX source in NGC 1097, designated ULX-3. To achieve this, we utilize archival data from multiple wavelengths, including X-ray observations from Chandra, XMM-Newton, and Swift/XRT, as well as optical data from the Hubble Space Telescope (HST). The precise positions of the ULX sources studied in both galaxies are shown on the Chandra images in Figure 1. Moreover, infrared data from the James Webb Space Telescope (JWST) are analyzed for the ULX in NGC 1097. The paper is organized as follows: Section 2 provides details of the observations. Data reduction and analysis of X-ray and optical observations are given in Section 3. Section 4 presents the results and discusses the properties of the ULXs. Finally, Section 5 summarizes the main findings of this study.
@@ 2 Observations
The NGC 4631 galaxy was observed by the Chandra ACIS-S instrument over 12 observations between 2000 and 2023. The galaxy also has XMM-Newton observations from 2002, two observations in 2021, and one observation in 2025. The galaxy was observed by the HST in 2003 and 2014 using the ACS/WFC (Advanced Camera for Surveys/Wide Field Channel). NGC 4631 was also observed 24 times by Swift/XRT between 2013 and 2021, with individual exposure times ranging from 128 s to 7369 s. In addition, between June 8, 2024, and January 9, 2025, the galaxy NGC 4631 was observed with the JWST/Mid-Infrared Instrument (MIRI) using the F1130W, F1500W, F1800W, F560W, and F770W filters. The JWST data were not used in this study because their spatial resolution (0.11 arcsec/pixel 4 pc/pixel) was insufficient to resolve the counterparts of X-6 and X-7.
NGC 1097 was observed with the Chandra ACIS-S instrument in six separate sessions between 2001 and 2025. The initial observation took place on January 28, 2001, followed by five recent observations in 2025—three on May 21 and two on May 31. The galaxy also has XMM-Newton data from an observation conducted on December 22, 2023. In addition, NGC 1097 was observed with JWST/NIRCam and MIRI using various filters between December 27, 2023, and September 29, 2024, and the galaxy has also been observed on multiple occasions with HST using different optical filters between June 5, 2004, and August 20, 2025. For Swift/XRT, 86 Photon Counting (PC) mode observations associated with Target IDs 00045597 and 00036582 were carried out between 2007 and 2025. More details of the observations of NGC 4631 and NGC 1097 are provided in Table 6.
3 X-ray and multiwavelength data reductions and analysis
@@ 3.1 X-ray energy spectra
The Chandra ACIS-S datasets were processed using the Chandra Interactive Analysis of Observations (CIAO) (Fruscione et al., 2006), along with calibration files from CALDB (v4.11). Level 2 event files were generated using the chandra_repro script available in CIAO. X-ray sources in the Chandra ACIS-S image were identified using the wavdetect tool from the CIAO. Wavelet scales of 2, 4, 8, and 16 pixels were applied, with a detection threshold set to 10-6. The positions of the sources newly classified as ULXs in this study, based on Chandra observations, are as follows: NGC 4631 X-6 at , ; NGC 4631 X-7 at , ; and NGC 1097 ULX-3 at , .
@@
Spectral data and light curves were derived through the specextract and dmextract tasks, respectively. For Chandra observations, we extracted the source counts using a circular region with a radius of 4′′. The background was estimated using two methods: (i) a nearby circular region with a radius of 8′′, free of point sources, and (ii) an annulus centered on the ULX with inner and outer radii of 6′′ and 10′′, respectively. Both approaches yielded consistent net count rates within the uncertainties. XMM-Newton EPIC observations were analyzed using the Science Analysis System (sas)111https://www.cosmos.esa.int/web/xmm-newton/sas v22.0. EPIC-PN and MOS event files were obtained with the rmfgen and arfgen tasks, respectively. The event selection criteria included PATTERN12 for EPIC-MOS and PATTERN4 for the PN, with FLAG=0 applied in both cases. Circular regions with radii of 15′′ and 30′′ were used for the extraction of source and background events, respectively, employing the evselect task. Since the ULXs NGC 4631 X-6 and NGC 1097 ULX-3 fall within the chip gap of the PN detector on XMM-Newton, only data from the MOS detectors were used in the analysis. The source energy spectra were grouped using the FTOOLS grppha. For the XMM-Newton MOS data, the spectra were grouped with at least 15-20 counts per energy bin, allowing the use of the statistic for the spectral fitting. In case of Chandra observations, the spectra were grouped with at least 5 and 10 counts per energy bin, and the C-statistic was used, as it is more appropriate for Poisson-distributed data. All spectral fits were performed within the 0.5–8 (for Chandra) and 0.3-10 keV (for XMM-Newton) energy range using XSPEC (Arnaud, 1996) to determine the best-fitting model for all spectra. In addition, because of the limited photon statistics in some Chandra observations and considering variability timescales, we combined spectra from observations with closely spaced dates to improve the statistical quality of the data and obtain the best spectral fits.
@@
For ULXs in NGC 4631, two-component models did not provide a statistically significant improvement over single-component models. Therefore, we adopted single-component models, either a power-law or a diskbb model combined with two absorption components (tbabs). We used two tbabs components: one fixed at the Galactic hydrogen column density (NH = 0.013 1022 cm-2) to account for Galactic absorption (Dickey and Lockman, 1990), and the other free to model intrinsic absorption. Intrinsic values were initially derived from the spectral fits, and the observed variations were considered more likely statistical than physical. To avoid these fluctuations mimicking flux changes, the fits were repeated with intrinsic fixed to the average of the initially obtained values. Intrinsic absorption variability in ULXs is, however, debated in the literature, with reports of both variable (e.g. Kajava and Poutanen, 2009) and more stable absorption properties (e.g. Gúrpide et al., 2021a). For NGC 1097 ULX-3, spectral modeling was performed only with the XMM-Newton observation because the Chandra observations had insufficient photon counts. A variety of standard spectral models were tested using MOS data, and the spectrum was best fit by a combined power-law + bbody model. The detailed model parameters for ULXs, along with the absorbed flux estimated using the cflux convolution model in XSPEC, are provided in Tables 1, 2, and 3 for X-6, X-7, and ULX-3, respectively. The energy spectra are displayed in Figures 2 and 3.
@@
Data label Net Rate NH /Tin /T (10-3counts/s) ( ) (/keV) (/keV) ( erg ) ( erg ) ( erg ) ( erg ) C1 ¡0.08 tbabs*tbabs*power-law C2 17.64 21.56/28b C3 7.93 27.48/29 C4 9.75 21.99/30 Ccombine1 9.06 58.47/60 C5 8.25 35.92/29 C6 9.63 32.27/22 C7 9.69 20.52/23 C8 4.59 32.46/23 Ccombine2 2.49 22.76/26 tbabs*tbabs*diskbb C2 17.62/28b C3 27.22/29 C4 23.44/30 Ccombine1 59/60 C5 36.04/29 C6 31.76/22 C7 17.82/23 C8 30.64/23 Ccombine2 20.85/26 XM1 ¡2.74 XM2 21.57 0.93 0.71 1.40 34.53/33 3.80 - - XM3 ¡2.11
@@
Notes: The of one tbabs component was fixed to the line-of-sight value of . The intrinsic (second tbabs component) was left to free to vary; additionally, for the entries marked with (*), the intrinsic was fixed at average values of and for the power-law and diskbb models, respectively. Ccombine1 and Ccombine2 correspond to combined observations C3–C4 and C9–C12, respectively. The count rate values for C9–C12 are (1.680.38)10-3, (2.070.32)10-3, (2.420.35)10-3, and (2.570.4410-3, respectively. The superscript ‘b’ indicates that the corresponding spectral fits were performed using the statistic. Absorbed fluxes were calculated in the 0.5–8 keV energy range, while unabsorbed luminosities were derived assuming a distance of 7.6 Mpc. Quoted uncertainties correspond to the 90% confidence level for each parameter.

### v2

# Identification of three new ultraluminous X-ray sources in NGC 4631 and NGC 1097: Evidence for stellar-mass black holes

@@ Abstract
Context.
Aims. Recent observations of galaxies continue to reveal new ultraluminous X-ray sources (ULXs), thereby increasing their known population. This growing sample provides improved statistics that are essential for advancing our understanding of ULX properties and their nature.
Methods. Our objective was to study the source populations using published Chandra, XMM-Newton, and Swift/XRT observations to identify new ULXs. In particular, we focused on investigating their X-ray and optical properties to constrain the nature of the compact objects and donor stars.
Results. We analyzed archival X-ray observations spanning 2000 to 2025 for NGC 4631 and NGC 1097 to study their ULX populations. We performed spectral fitting for sources with sufficient counts using absorbed power-law and diskbb models to determine their physical properties. We conducted variability analyses, including hardness-intensity diagrams and light curves, to assess both short- and long-term variability. Additionally, we used optical color-magnitude diagrams and near-infrared (NIR) spectral energy distributions (SEDs) to identify and characterize possible donor stars.
Conclusions. In NGC 4631, we identify two new transient ULXs, X-6 and X-7, whose X-ray count rates vary by more than an order of magnitude. The relation obtained from the diskbb fits provides strong evidence that NGC 4631 X-6 is powered by a stellar-mass black hole (BH) accreting via a standard disk. The optical sources within the X-6 and X-7 X-ray error circles are candidate optical counterparts, suggesting that these systems are candidate high-mass X-ray binaries. In NGC 1097, we report the discovery of a new transient ULX, designated ULX-3, which exhibits variations in X-ray luminosity (LX) by a factor of about 30. For ULX-3, we detect spectral state transitions, which may indicate a compact object consistent with either a stellar-mass BH or a neutron star. Moreover, we identify a unique optical and NIR counterpart: while the optical emission is variable, the NIR emission remains stable. The NIR counterpart of ULX-3 shows an SED consistent with a blackbody temperature of K, which lies within the expected range for red supergiants with a stellar radius of .
@@
Ultraluminous X-ray sources (ULXs) are the most extreme subclass of X-ray binaries (XRBs), appearing as off-nuclear point sources exhibiting ¿ 2 erg s-1 (Kaaret et al., 2017; Fabrika et al., 2021; King et al., 2023; Pinto and Walton, 2023). To account for their ultraluminosities, recent studies increasingly support the view that the majority of ULXs are powered by stellar-mass compact objects undergoing super-Eddington accretion (King et al. (2023) and references therein), rather than by sub-Eddington accretion onto intermediate-mass black holes(Heinzeller et al., 2006; Brightman et al., 2019; Greene, 2022). Within this framework, the accreting compact object is generally inferred to be either a black hole (BH) or a neutron star (NS).
@@
Coherent pulsations have been detected in several ULXs, confirming them as pulsating NS ULXs (PULXs), primarily through observations by XMM-Newton and NuSTAR (Bachetti et al., 2014; Motch et al., 2014; Israel et al., 2017; Carpano et al., 2018). Moreover, the observation of cyclotron resonance scattering features in the X-ray spectrum of a ULX implies the presence of an NS candidate (Mushtukov et al., 2017; Pinto et al., 2020). In addition, several ULXs, including some identified as PULXs, show transient behavior, whereas most remain persistently bright in the X-ray band over long timescales. Repeated observations of nearby galaxies with Chandra and XMM-Newton have contributed to a growing number of ULXs, particularly those exhibiting transient behavior (van Haaften et al., 2019; Earnshaw et al., 2020; Walton et al., 2021; Robba et al., 2022; Roberts et al., 2023; Brightman et al., 2023; Allak, 2023).
@@
Several ULXs also show clear evidence of recurrent spectral state transitions, evolving from hard to soft regimes over time, as revealed by high-cadence and long-term studies. In particular, Holmberg II X1 and NGC 5204 X1 show a cyclical evolution through the states of hard ultra-luminous (HUL), soft ultra-luminous (SUL), and supersoft ultra-luminous (SSUL), likely driven by variations in accretion rate and changes in funnel geometry caused by radiation‑driven winds (Gúrpide et al., 2021b; Mondal et al., 2021). A recent study by Majumder et al. (2025) analyzed the long-term evolution of several ULXs with candidate BH accretors using XMM-Newton, revealing diverse spectrotemporal behaviors and correlations between disk luminosity (Ldisc) and inner-disk temperature (Tcol). The authors argue that these correlations are indicative of possible spectral states or transitions between standard thin-disk and slim-disk configurations at different luminosities, supporting a two-component supercritical accretion framework regulated by the disk, corona, and radiation-driven winds. Numerous studies have sought to identify the optical counterparts of ULXs and clarify the origin of their emission (Tao et al., 2011; Soria et al., 2012; Gladstone et al., 2013; Vinokurov et al., 2018; Allak et al., 2022; Guo et al., 2023; Allak and Akyuz, 2025). However, determining the precise contributions remains challenging, as the emission may originate from the donor star’s photosphere, the accretion disk, or a combination of both (Tao et al., 2011; Fabrika et al., 2015; Allak et al., 2022). In rare cases, spectral signatures such as stellar absorption lines allow clear donor identification, as seen in NGC 7793 P13, where a B9Ia supergiant was confirmed (Motch et al., 2014). Optical variability has also been observed in several ULXs, such as ULX-4 and ULX-8 (both in M51) and NGC 1313 X-2, suggesting accretion disc-dominated emission in these systems (Liu et al., 2009; Allak, 2022). Optical counterparts of ULXs are generally faint (mV ¿ 20 mag) and frequently obscured by surrounding stellar populations or interstellar material, which complicates their detection and interpretation. In contrast, infrared (IR) observations, particularly in the NIR range, are less affected by dust extinction and provide valuable insights into the nature of donor stars and the circumstellar environments of ULXs. Several ULXs have reported NIR magnitudes consistent with red supergiants (RSGs) Heida et al., 2014; Lau et al., 2019; Dudik et al., 2016, and spectroscopic analyses suggest RSG-like donors in some cases (Heida et al., 2016). Mid-IR emission is also attributed to either circumstellar dust or jets (Dudik et al., 2016; Lau et al., 2019; Allak, 2023). NGC 4631, also known as the Whale Galaxy, is a barred spiral galaxy that exhibits starburst activity. It is located in the constellation Canes Venatici, approximately 7.6 Mpc away (1′′ = 34 pc) (Soria and Ghosh, 2009). Several studies have explored the ULX population in NGC 4631, revealing a range of accretion processes. Carpano et al. (2007) reported the discovery of a supersoft ULX that exhibits a 4-hour periodic modulation based on XMM-Newton data. This supersoft ULX is unlikely to be explained by the standard white dwarf scenario without extreme beaming. It has been suggested that it may instead be a stellar-mass BH accreting at super-Eddington rates, with the modulation resulting from either a companion eclipse or a warped accretion disk.
@@
Furthermore, using archival data from XMM-Newton, Chandra, and ROSAT, Soria and Ghosh (2009) reanalyzed the five brightest X-ray sources in NGC 4631, identifying four (X1, X2, X4, and X5) as ULXs. Among these, X1—previously described as a supersoft source by Carpano et al. (2007)—has been proposed to represent an extreme subclass of supersoft sources, powered by transient super-Eddington outbursts driven by non-steady nuclear burning on the surface of a massive white dwarf. The other sources displayed spectral properties consistent with stellar-mass BHs accreting at super-Eddington rates. Moreover, Ducci et al. (2025) report the discovery of a new transient pulsar ULX in NGC 4631, designated X–8, with a spin period of and a very high spin-up rate, . Guo et al. (2023) studied the optical counterparts of ULXs in NGC 4631 using a broadband and narrowband imaging campaign with Canada–France–Hawaii Telescope (CFHT)/MegaCamm, aiming to detect bubble structures around the X-ray sources and examining their accretion properties. They analyzed the stellar surroundings of the sources via extinction-corrected color–magnitude diagrams and isochrone fitting. Their study reveals a highly asymmetric bubble nebula around X4, displaying distinct morphologies in H and [O III] emission.
@@
NGC 1097 (D = 16.8 Mpc, 1′′ = 82 pc; Liu 2011) is typically classified as a barred spiral (SBb) galaxy. It is also classified as a Seyfert 2 galaxy, which hosts an active galactic nucleus (AGN) powered by a central supermassive BH. Consequently, most studies of this galaxy have focused on constructing the SED of its nucleus using X-ray and multiwavelength observations (Nemmen et al., 2006). Observations with Chandra data further revealed two ULX sources, X2 and X3, in the galaxy, with model-estimated luminosities of erg s-1 and erg s-1, respectively (Liu, 2011).
@@
In this study, we investigate the nature of two newly identified ULX sources in NGC 4631, designated X-6 and X-7, and one ULX source in NGC 1097, designated ULX-3. To achieve this, we used archival data from multiple wavelengths, including X-ray observations from Chandra, XMM-Newton, and Swift/XRT, as well as optical data from the Hubble Space Telescope (HST). The precise positions of the ULX sources studied in both galaxies are shown on the Chandra images in Fig. 1. In addition, we analyzed infrared data from the James Webb Space Telescope (JWST) for the ULX in NGC 1097. The paper is organized as follows: Section 2 provides details of the observations. Section 3 describes the data reduction and analysis of X-ray and optical observations. Section 4 presents the results and discusses the properties of the ULXs. Finally, Section 5 summarizes the main findings of this study.
@@ 2 Observations
The NGC 4631 galaxy was observed with the Chandra ACIS-S instrument over 12 observations between 2000 and 2023. Archival observations of the galaxy also include data from XMM-Newton taken in 2002, two observations in 2021, and one in 2025. The HST observed the galaxy in 2003 and 2014 using the Advanced Camera for Surveys/Wide Field Channel (ACS/WFC). NGC 4631 was also observed 24 times by Swift/XRT between 2013 and 2021, with individual exposure times ranging from 128 to 7369 s. In addition, between 8 June 2024 and 9 January 2025, the JWST/Mid-Infrared Instrument (MIRI) observed NGC 4631 using the F1130W, F1500W, F1800W, F560W, and F770W filters. We did not use the JWST data in this study because their spatial resolution (0.11 arcsec/pixel 4 pc/pixel) was insufficient to resolve the counterparts of X-6 and X-7.
The Chandra ACIS-S instrument observed NGC 1097 in six separate sessions between 2001 and 2025. The initial observation took place on January 28, 2001, followed by five observations in 2025 (three on 21 May, two on 31 May). The galaxy also has XMM-Newton data from an observation conducted on 22 December 2023. In addition, the JWST/NIRCam (Near-Infrared Camera) and MIRI (Mid-Infrared Instrument) instruments observed NGC 1097 with multiple filters between 27 December 2023 and 29 September 2024. The HST also observed the galaxy on multiple occasions using different optical filters between 5 June 2004 and 20 August 2025. For Swift/XRT, 86 photon counting (PC) mode observations associated with Target IDs 00045597 and 00036582 were carried out between 2007 and 2025. Table 6 provides additional details on the observations of NGC 4631 and NGC 1097.
3 X-ray and multiwavelength data reduction and analysis
@@ 3.1 X-ray energy spectra
We processed the Chandra ACIS-S datasets using Chandra Interactive Analysis of Observations (CIAO) (Fruscione et al., 2006) and calibration files from the Calibration Database (CALDB) version 4.11, generating level 2 event files using the chandra_repro script available in CIAO. Using the wavdetect tool in CIAO, we identified X-ray sources and applied wavelet scales of 2, 4, 8, and 16 pixels with a detection threshold set to 10-6. The positions of the sources newly classified as ULXs in this study, based on Chandra observations, are: NGC 4631 X-6 at , ; NGC 4631 X-7 at , ; and NGC 1097 ULX-3 at , .
@@
We derived spectral data and light curves with specextract and dmextract, respectively. For Chandra observations, we extracted the source counts using a circular region with a radius of 4′′. We used two methods to estimate the background: (i) a nearby circular region with a radius of 8′′, free of point sources, and (ii) an annulus centered on the ULX with inner and outer radii of 6′′ and 10′′, respectively. Both approaches yield consistent net count rates within uncertainties. We analyzed XMM-Newton European Photon Imaging Camera (EPIC) observations using the science analysis system (sas)111https://www.cosmos.esa.int/web/xmm-newton/sas v22.0. Response files (RMFs and ARFs) for the EPIC-pn and MOS (Metal Oxide Semiconductor detectors) were generated using the rmfgen and arfgen tasks, respectively. The event selection criteria included PATTERN12 for European Photon Imaging Camera – Metal Oxide Semiconductor (EPIC-MOS) and PATTERN4 for PN, with FLAG=0 applied in both cases. We extracted source and background events from circular regions with radii of 15′′ and 30′′, respectively, using the evselect task. Because NGC 4631 X-6 and NGC 1097 ULX-3 fall into the chip gap of the PN detector on XMM-Newton, we used data from the MOS detectors only. We used FTOOLS grppha to group the source energy spectra. For the XMM-Newton MOS data, we grouped the spectra with at least 15-20 counts per energy bin to enable spectral fitting. For Chandra observations, the spectra were grouped with 5-10 counts per energy bin, using C-statistics suitable for Poisson-distributed data. We performed all spectral fits using XSPEC (Arnaud, 1996), covering 0.5–8 keV (Chandra) and 0.3-10 keV (XMM-Newton), to determine the best-fitting models. In addition, due to limited photon statistics in some Chandra observations and considering variability timescales, we combined spectra from closely spaced observations to improve the data’s statistical quality and obtain the best spectral fits.
@@
For NGC 4631 ULXs, two-component models did not provide a statistically significant improvement over single-component models. Therefore, we adopted single-component models, either a power-law or a diskbb model combined with two absorption components (tbabs). We used two tbabs components: one fixed at the Galactic hydrogen column density (NH = 0.013 1022 cm-2) to account for Galactic absorption (Dickey and Lockman, 1990), and the other free to model intrinsic absorption. We initially derived intrinsic values from the spectral fits and interpreted the observed variations as more likely statistical than physical. To avoid these fluctuations mimicking flux changes, we repeated the fits with intrinsic fixed to the average of the initial values obtained. Intrinsic absorption variability in ULXs remains debated, with reports of both variable (e.g., Kajava and Poutanen, 2009) and more stable absorption properties (e.g., Gúrpide et al., 2021a). For NGC 1097 ULX-3, we performed spectral modeling using only the XMM-Newton observation, because the Chandra observations had insufficient photon counts. with the spectrum best fit by a combined power-law plus bbody model. We provide detailed model parameters for the ULXs, along with the absorbed flux estimated using the cflux convolution model in XSPEC, in Tables 1, 2, and 3 for X-6, X-7, and ULX-3, respectively. Figures 2 and 3 show the corresponding energy spectra.
@@
Data label Net Rate NH /Tin /T (10-3counts/s) ( ) (/keV) (/keV) ( erg ) ( erg ) ( erg ) ( erg ) C1 ¿0.08 tbabs*tbabs*power-law C2 17.64 21.56/28b C3 7.93 27.48/29 C4 9.75 21.99/30 Ccombine1 9.06 58.47/60 C5 8.25 35.92/29 C6 9.63 32.27/22 C7 9.69 20.52/23 C8 4.59 32.46/23 Ccombine2 2.49 22.76/26 tbabs*tbabs*diskbb C2 17.62/28b C3 27.22/29 C4 23.44/30 Ccombine1 59/60 C5 36.04/29 C6 31.76/22 C7 17.82/23 C8 30.64/23 Ccombine2 20.85/26 XM1 ¿2.74 XM2 21.57 0.93 0.71 1.40 34.53/33 3.80 - - XM3 ¿2.11
@@
Notes: The of one tbabs component was fixed to the line-of-sight value of . The intrinsic (second tbabs component) was left to free to vary; for the entries marked with (*), the intrinsic was fixed at average values of and for the power-law and diskbb models, respectively. Ccombine1 and Ccombine2 correspond to combined observations C3–C4 and C9–C12, respectively. The count rate values for C9–C12 are (1.680.38)10-3, (2.070.32)10-3, (2.420.35)10-3, and (2.570.4410-3, respectively. The superscript b indicates that the corresponding spectral fits were performed using the statistic. Absorbed fluxes were calculated in the 0.5–8 keV energy range, while unabsorbed luminosities were derived assuming a distance of 7.6 Mpc. Quoted uncertainties correspond to the 90% confidence level for each parameter